OpenStudy (anonymous):

If m∠1+m∠2=90 and m∠2=m∠3 then m∠1=m∠3=90 What property is this

OpenStudy (anonymous):

One that doesn't make any sense. Did you copy the last equation down correctly?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

Did you copy *all* the equations down correctly? Here's a proof of why the question doesn't make sense: If m∠1+m∠2=90 and m∠2=m∠3 Then m∠3+m∠2=90 If m∠2=m∠3 and m∠3+m∠2=90 Then m∠3+m∠3=90 If m∠3=90 (see Given III) and m∠3+m∠3=90 Then 90+90=90 ...Which is false.

OpenStudy (anonymous):

m∠1+m∠2=90 and m∠2=m∠3 then, m∠1+m∠3=90

OpenStudy (anonymous):

Im asking for the Property as in transitive or reflexive so on and so on